### v2

@@ Introduction
For millennia, human-created content served as the exclusive record of civilization, with generally every painting, text, and diagram traceable to human thought and effort [1, 2, 3]. Generative AI has brought an end to this era [4], producing seemingly indistinguishable alternatives at unprecedented scale [5, 6, 7], permeating every domain from scientific publications [8, 9] to medical diagnostics [10, 11] to artistic designs [12, 13]. Yet this transformative technology raises profound concerns about its deployment across critical domains: from diagnostic reliability in medical applications [14, 15, 16] and accelerated misinformation propagation [17, 18] to cultural homogenization [19, 20, 21] and data contamination [22, 23]. These consequential risks necessitate immediate development of rigorous methods to analyze, evaluate, and control AI-generated content, protecting the integrity and trustworthiness of AI systems in high-stakes applications.
@@
Content analysis for generative AI encounters a fundamental scope-specificity dilemma. While effective content analysis requires comprehensive assessment across multiple quality dimensions, from prompt adherence to anatomical correctness, real-world failures typically manifest as specific visual patterns that require precisely the granular detection capabilities that many broad-scope methods fail to provide [24]. Existing analytic methods typically prioritize either comprehensive scope or detection specificity. Some analytic approaches treat generated content as indivisible units, providing holistic assessments through distributional alignment metrics [25, 26] or semantic similarity scores [27, 28], while other developments invent specialized detection tools for specific failure modes—such as text generation errors [29, 30], model hallucination [31], or anatomical anomalies [32]. These specialized approaches achieve high precision but lack adaptability across different content domains.
Comparatively, human evaluation of AI-generated content operates through fundamentally different mechanisms than automated analytic approaches. Consider how art critics assess AI-generated paintings: rather than providing holistic scores, they articulate which aspects work and which fail: the lighting may be striking, the composition balanced, the anatomy flawed, which are all inherently visual patterns interpreted with natural language. This pattern-based evaluation process reflects broader cognitive mechanisms documented in human visual assessment [33, 34, 35, 36, 37, 38, 39, 40], suggesting a promising approach to content analysis: developing methods that identify language-grounded visual patterns in AI-generated content.
Here we introduce Language-Grounded Sparse Encoders (LanSE), a content analysis method that systematically identifies interpretable visual patterns in AI-generated content and grounds them with natural language descriptions. LanSE leverages recent advances in mechanistic interpretability [41, 42], particularly sparse autoencoders [43, 44, 45] and transcoders [46, 47, 48], combined with large multimodal models (LMM) analysis to comprehensively discover, interpret and organize visual patterns across multiple content aspects. We validate LanSE’s content analysis capability through extensive human evaluation involving over 11,000 annotations from six independent annotators. This evaluation demonstrates strong alignment with human judgment, achieving 93% agreement when detecting visual patterns in natural images. This robust validation establishes LanSE as a reliable foundation for systematic content analysis across diverse applications.
LanSE’s core capability to comprehensively identify visual patterns enables three complementary applications in content analysis for generative AI. For generative model evaluation, LanSE enables systematic assessment through four evaluation metrics: prompt match (whether concepts in prompts exist in the output [49]), visual realism (whether the output is photorealistic [50]), physical plausibility (whether the output adheres to physics constraints [51]), and content diversity (whether outputs show sufficient variation [52]). LanSE enables systematic automated evaluation of physical plausibility in AI-generated content and reveals failure modes that existing methods overlook. For content understanding, LanSE provides interpretable multi-label annotations that enable natural language-based image retrieval, systematic dataset curation, and granular failure mode detection across thousands of visual patterns. Beyond single-model analysis, LanSE enables inter-model correlation analysis that reveals how different generative systems encode visual knowledge, showing that models converge on semantic representations while diverging in their specific failure modes. For cross-domain analysis, the core methodology extends naturally beyond natural images: we demonstrate clinical pattern detection in medical images achieving 74% agreement with expert radiologists, with potential applications spanning satellite imagery analysis, biological microscopy analysis, and scientific data interpretation. We believe LanSE represents a critical step toward transparent and accountable AI systems, addresses the content analysis need for this generative AI era, and paves the way to interpretable AI architectures where human understanding guides system behavior.
@@ Results
LanSE systematically discovers visual patterns by collecting millions of candidate neurons from interpretability modules, using LMMs to describe the visual patterns each neuron represents, and retaining only those with accuracies greater than 80% (detailed in Method). This process yields 5,309 high-quality neurons, each corresponding to a distinct visual pattern found in AI-generated images. These patterns are grouped into three main categories with nine subgroups: semantic content—(1) human, (2) animal, (3) object, (4) activity, (5) environment; unrealistic patterns—(6) style, (7) artifact; and physics violations—(8) distortion, (9) structure. Each category plays a specific role in LanSE’s evaluation method (examples in Appendix A).
To understand if the visual patterns detected by LanSE are both accurate and reliable for evaluating image quality, we validated them against large-scale human and automated assessments. We conducted evaluations involving over 11,160 human annotations from four independent annotators (see Appendix F). Each human annotator reviewed 2,790 neuron–image pairs, where each pair comprised a neuron’s interpreted description and an image that activated it, and judged whether the activation matched the described pattern. As shown in Figure 2, visual patterns discovered by LanSE achieved 93.2% accuracy under human evaluation, with category-specific performance ranging from 86.0% to 99.9%. The four annotators achieved an inter-annotator agreement of 85.93%. In addition to human annotation, we also included automated assessments from two state-of-the-art LMMs. On the same dataset, automated evaluation via LMM shows strong agreement with human annotation achieving 90.4% accuracy with category-specific performance ranging from 76.2% to 98.5%.
@@
LanSE provides detailed multi-label annotations for individual images by identifying all activated visual patterns, enabling comprehensive content characterization at the datapoint level as demonstrated in Figure 3. Across our test set of 48,000 generated images, we observe that images typically activate 15-30 of the 5,309 LanSE neurons. Individual images reveal complex pattern combinations—for instance, simultaneously activating neurons for ”depiction of human figures” (human category), ”transportation-related elements in urban settings” (environment category), and ”anatomically impossible structures” (structure category), providing comprehensive characterization of image samples.
@@
LanSE’s comprehensive visual pattern vocabulary enables the construction of four diagnostic metrics that systematically characterize different aspects of generated content: prompt match, visual realism, physical plausibility, and content diversity (detailed definition in Appendix E). These metrics target complementary aspects that various existing metrics fail to distinguish as shown in Figure 4.
@@
To demonstrate LanSE’s advantages for content analysis, we compare our diagnostic capabilities with widely-adopted baseline methods: Fréchet Inception Distance (FID) [25], Inception Score (IS) [26], and CLIP score [27].
Our analysis aligns with CLIP’s systematic biases identified in various previous works [67]. While both CLIP and our prompt match metric generally detect severe misalignments (empty upper-right regions in Figure 10), CLIP can assign high scores to outputs with obvious semantic mismatches when partial object alignment masks compositional errors, as shown in Figure 9. In contrast, LanSE’s prompt match evaluation is based on visual patterns and can capture subtle semantic mismatches CLIP fails to detect.
@@
In this era of generative AI, many researchers consider LMMs with prompt engineering as a quick and universal solution for diverse problems [68, 69, 70]. Comparatively, LanSE can detect fine-grained visual patterns that state-of-the-art LMMs miss, while requiring significantly lower computational resources during inference. We compare LanSE’s specialized error detection capabilities against six state-of-the-art vision-language models on identifying physical plausibility errors. Using 3,410 human-annotated images (2,522 with physics violations, 888 without), we tasked each system with detecting anatomical impossibilities, distorted structures, and physics violations.
@@
We are witnessing a profound societal transformation as AI-generated content gradually permeates nearly every domain of human content creation, fundamentally challenging how we understand, verify, and trust the materials we encounter. Research across multiple domains documents serious societal risks from AI-generated content proliferation, including homogenization of human expression [20] and creative diversity [19], vulnerability to data poisoning attacks in medical systems [14], model collapse from recursive training on synthetic data [22], contamination of scientific datasets with undetected artificial material [71], and erosion of editorial standards in professional publishing [17]. These documented challenges underscore the critical need for comprehensive content analysis infrastructure that can systematically characterize AI-generated content properties, enabling informed governance and deployment of these materials across high-stakes domains. LanSE provides the systematic content analysis infrastructure this transformed landscape demands, automatically discovering thousands of interpretable visual patterns that enable comprehensive, human-understandable assessment of AI-generated content properties.
@@
LanSE’s systematic pattern analysis transforms how critical institutions—from hospitals to regulatory agencies—understand and govern AI-generated content. In medical diagnostics and financial decision-making, LanSE provides clinicians and analysts with language-grounded explanations of AI predictions, delivering the transparency and accountability these domains require. Regulators can customize LanSE’s pipeline to detect domain-specific risks for educational, publishing, and safety-sensitive materials, enabling precise content filtering with transparent natural language justifications. For machine learning researchers, LanSE enables systematic dataset analysis by quantifying pattern distributions to reveal bias and guide targeted augmentation, while providing early warning signals for model collapse through diversity measurement. Beyond analysis applications, LanSE enables integration into training pipelines for interpretable model feedback and supports development of pattern-based retrieval systems where users can find content sharing specific visual characteristics. For the interpretability field, LanSE represents a paradigm shift—repositioning discovered patterns from primarily explanatory tools to functional infrastructure for diverse downstream applications.
While LanSE demonstrates versatility across multiple applications, several avenues could strengthen and extend this method. First, expanding the scale and granularity of discovered visual patterns would enhance evaluation precision, as current assessments remain bounded by our pattern vocabulary. Second, although we successfully adapted LanSE to medical imaging through CXR-LanSE, extending the framework to additional modalities, video, 3D content, or scientific imaging, presents promising opportunities. Third, while we currently demonstrate LanSE primarily for content analysis, dataset characterization, and model correlation studies, we envision integrating LanSE directly into model training pipelines, enabling generative models to receive diverse, interpretable feedback on specific visual patterns during optimization. Fourth, developing adaptive pattern vocabularies that evolve with emerging AI capabilities could maintain LanSE’s effectiveness as generative models advance.
@@
We use a collection of image-caption pairs from MS-COCO [72], Flickr8k and Flickr30k [73], TextCaps [74], NoCaps [75], and MIMIC-CXR [76] to build LanSE and CXR-LanSE. We augment these datasets with samples from several generative models [62, 63, 64, 65, 77, 78] by prompting them with captions from the original collections. We also employ human data annotators to identify erroneous images from the dataset for targeted discovery of physical plausibility patterns (detailed in Appendix D).
@@
To systematically discover interpretable visual patterns for content analysis, we leverage recent advances in mechanistic interpretability [43, 44, 45, 46, 47, 48]. Studies have shown that enforcing sparsity during neural network training produces monosemantic neurons—individual units that respond to specific, interpretable concepts [79, 80, 81, 82, 83], motivating us to adapt these techniques for content analysis.
@@
Sparse autoencoders are sparsity-driven interpretability modules that can comprehensively discover a diverse set of general visual patterns through the reconstruction of one model embedding. We develop multiple sparse autoencoders, each comprising a single-layer encoder $\mathbf{z_{i}}=\mathrm{Enc}(\mathbf{e_{i}})=\mathrm{ReLU}(\mathbf{w_{enc}}\cdot\mathbf{e_{i}}+\mathbf{b_{enc}})$ and a single-layer decoder $\mathrm{Dec}(\mathbf{z_{i}})=\mathrm{ReLU}(\mathbf{w_{dec}}\cdot\mathbf{z_{i}}+\mathbf{b_{dec}})$. A top-$k$ activation scheme [45] is used to encourage sparsity in the latent space: for each $\mathbf{z_{i}}$, we retain only the $k$ highest-magnitude activations and set the rest to zero. The training objective is as follows:
@@
The neurons discovered by these interpretability modules are initially raw and noisy, requiring systematic curation to identify reliably interpretable patterns. To maximize pattern discovery coverage while enabling quality filtering, we train multiple modules on different data subsets. In total, we train 200 sparse autoencoders and 50 transcoders for LanSE, plus 100 sparse autoencoders for CXR-LanSE, with each module containing 15,000 latent neurons and k=32 for the top-k sparsity constraint. From this large pool of candidate neurons, we apply systematic filtering criteria to identify those with consistent, interpretable activations. We adopt hyperparameters similar to recent developments in mechanistic interpretability [84, 85] and the approach can be scaled with additional compute and data.
@@
After training the interpretability modules, we obtain a collection of latent neurons, each of them potentially interpretable in natural language. We then employ LMMs [86, 87] to analyze and interpret these latent neurons.
For each neuron, we collect a set of samples that trigger strong activations for this neuron and prompt Claude-3.5-Haiku [86] (a large multimodal model) to summarize the common characteristics within these samples. The prompt we use to summarize the visual pattern in natural language is given below (more prompts in Appendix C):
@@
Based on the natural language summarizations of the visual patterns, we categorize these visual patterns into nine groups, $\mathbf{C}$, including five groups designed for semantics: human, animal, object, activity, and environment; two groups for visual realism: style and artifact; and two groups for physical plausibility: distortion and structure (See Appendix G). We use the following prompt to categorize the visual pattern based on their explanations.
@@
These visual pattern analysis procedures can be formulated as below [88], noticing that the summarizations are obtained automatically during the construction of LanSEs, no assistance from LMMs are used in inference.
@@
Modality distillation refers to training single-modality LanSEs, $\mathrm{S^{img}_{\text{c}}}$ and $\mathrm{S^{txt}_{\text{c}}}$, using outputs from $\mathrm{S_{\text{c}}}(\mathbf{x},\mathbf{y})$ [89]. The LanSEs are initialized with pretrained encoders and tuned with parameter efficient fine-tuneing [90]. The optimization target is to minimize the distances between the outputs of the single-modality LanSEs and $\mathrm{S_{\text{c}}}(\mathbf{x},\mathbf{y})$:

### v3

@@ Introduction
For millennia, human-created content served as the exclusive record of civilization, with generally every painting, text, and diagram traceable to human thought and effort [morriss2010evolution, raaflaub2013thinking, gripshover2020writing]. Generative AI has brought an end to this era [LivingBraveAI2023], producing seemingly indistinguishable alternatives at unprecedented scale [review2024generative, porter2024ai, nightingale2022ai], permeating every domain from scientific publications [gil2022will, kitano2021nobel, geng2024impactlargelanguagemodels] to medical diagnostics [mahmood2019deep, teixeira2018generatingsyntheticxrayimages] to artistic designs [rombach2022highresolutionimagesynthesislatent, kottapallitransforming]. Yet this transformative technology raises profound concerns about its deployment across critical domains: from diagnostic reliability in medical applications [alber2025medical, hager2024evaluation, kundu2023measuring] and accelerated misinformation propagation [lindsay2023llms, feng2025unravelingmisinformationpropagationllm] to cultural homogenization [doshi2024generative, suzuki2023we, choudhury2023generative] and data contamination [shumailov2024ai, smith2024ai]. These consequential risks necessitate immediate development of rigorous methods to analyze, evaluate, and control AI-generated content, protecting the integrity and trustworthiness of AI systems in high-stakes applications [10.1093/polsoc/puaf001, KRAKOWSKI2025100560].
@@
Content analysis for generative AI encounters a fundamental scope-specificity dilemma. While effective content analysis requires comprehensive assessment across multiple quality dimensions, from prompt adherence to anatomical correctness, real-world failures typically manifest as specific visual patterns that require precisely the granular detection capabilities that many broad-scope methods fail to provide [chen2023understandingvulnerabilityclipimage]. Existing analytic methods typically prioritize either comprehensive scope or detection specificity. Some analytic approaches treat generated content as indivisible units, providing holistic assessments through distributional alignment metrics [jayasumana2024rethinkingfidbetterevaluation, barratt2018noteinceptionscore] or semantic similarity scores [chen2025contrastivelocalizedlanguageimagepretraining, hessel2022clipscorereferencefreeevaluationmetric], while other developments invent specialized detection tools for specific failure modes—such as text generation errors [yao2024hifi, sampaio2024typescoretextfidelitymetric], model hallucination [li2023haluevallargescalehallucinationevaluation], or anatomical anomalies [fang2024humanrefinerbenchmarkingabnormalhuman]. These specialized approaches achieve high precision but lack adaptability across different content domains.
@@
Comparatively, human evaluation of AI-generated content operates through fundamentally different mechanisms than automated analytic approaches. Consider how art critics assess AI-generated paintings: rather than providing holistic scores, they articulate which aspects work and which fail: the lighting may be striking, the composition balanced, the anatomy flawed, which are all inherently visual patterns interpreted with natural language. This pattern-based evaluation process reflects broader cognitive mechanisms documented in human visual assessment [biederman1987recognition, zeki1999art, clark1996using, nightingale2017can, farid2022creating, rhodes1998facial, Santos03042018, landy2013texture], suggesting a promising approach to content analysis: developing methods that identify language-grounded visual patterns in AI-generated content.
Here we introduce Language-Grounded Sparse Encoders (LanSE), a content analysis method that systematically identifies interpretable visual patterns in AI-generated content and grounds them with natural language descriptions (See Figure 1). LanSE leverages recent advances in mechanistic interpretability [templeton2024scaling, lieberum2024gemmascopeopensparse], particularly sparse autoencoders [cunningham2023sparseautoencodershighlyinterpretable, bussmann2025learningmultilevelfeaturesmatryoshka, gao2024scalingevaluatingsparseautoencoders] and transcoders [dunefsky2024transcodersinterpretablellmfeature, paulo2025transcodersbeatsparseautoencoders, lindsey2024sparse], combined with large multimodal models (LMM) analysis to comprehensively discover, interpret and organize visual patterns across multiple content aspects. We validate LanSE’s content analysis capability through extensive human evaluation involving over 11,000 annotations from six independent annotators. This evaluation demonstrates strong alignment with human judgment, achieving 93% agreement when detecting visual patterns in natural images. This robust validation establishes LanSE as a reliable foundation for systematic content analysis across diverse applications.
LanSE’s core capability to comprehensively identify visual patterns enables three complementary applications in content analysis for generative AI. For generative model evaluation, LanSE enables systematic assessment through four evaluation metrics: prompt match (whether concepts in prompts exist in the output [huang2025t2icompbenchenhancedcomprehensivebenchmark]), visual realism (whether the output is photorealistic [kynkäänniemi2019improvedprecisionrecallmetric]), physical plausibility (whether the output adheres to physics constraints [yang2018physicsinformeddeepgenerativemodels]), and content diversity (whether outputs show sufficient variation [gerstgrasser2024modelcollapseinevitablebreaking]). LanSE enables systematic automated evaluation of physical plausibility in AI-generated content and reveals failure modes that existing methods overlook. For content understanding, LanSE provides interpretable multi-label annotations that enable natural language-based image retrieval, systematic dataset curation, and granular failure mode detection across thousands of visual patterns. Beyond single-model analysis, LanSE enables inter-model correlation analysis that reveals how different generative systems encode visual knowledge, showing that models converge on semantic representations while diverging in their specific failure modes. For cross-domain analysis, the core methodology extends naturally beyond natural images: we demonstrate clinical pattern detection in medical images achieving 74% agreement with expert radiologists, with potential applications spanning satellite imagery analysis, biological microscopy analysis, and scientific data interpretation. We believe LanSE represents a critical step toward transparent and accountable AI systems, addresses the content analysis need for this generative AI era, and paves the way to interpretable AI architectures where human understanding guides system behavior.
@@ Results
LanSE systematically discovers visual patterns by collecting millions of candidate neurons from interpretability modules, using LMMs to describe the visual patterns each neuron represents, and retaining only those with accuracies greater than 80% (detailed in Method). This process yields 5,309 high-quality neurons, each corresponding to a distinct visual pattern found in AI-generated images. These patterns are grouped into three main categories with nine subgroups: semantic content—(1) human, (2) animal, (3) object, (4) activity, (5) environment; unrealistic patterns—(6) style, (7) artifact; and physics violations—(8) distortion, (9) structure. Each category plays a specific role in LanSE’s evaluation method (examples in Appendix Appendix A: Qualitative Examples of LanSE Neurons).
To understand if the visual patterns detected by LanSE are both accurate and reliable for evaluating image quality, we validated them against large-scale human and automated assessments. We conducted evaluations involving over 11,160 human annotations from four independent annotators (See Appendix Appendix F. Human Evaluation Protocols). Each human annotator reviewed 2,790 neuron–image pairs, where each pair comprised a neuron’s interpreted description and an image that activated it, and judged whether the activation matched the described pattern. As shown in Figure 2, visual patterns discovered by LanSE achieved 93.2% accuracy under human evaluation, with category-specific performance ranging from 86.0% to 99.9%. The four annotators achieved an inter-annotator agreement of 85.93%. In addition to human annotation, we also included automated assessments from two state-of-the-art LMMs. On the same dataset, automated evaluation via LMM shows strong agreement with human annotation achieving 90.4% accuracy with category-specific performance ranging from 76.2% to 98.5%.
@@
LanSE provides detailed multi-label annotations for individual images by identifying all activated visual patterns, enabling comprehensive content characterization at the datapoint level as demonstrated in Figure 3. Across our test set of 48,000 generated images, we observe that images typically activate 15-30 of the 5,309 LanSE neurons (see Appendix Appendix H. Ablation Studies). Individual images reveal complex pattern combinations—for instance, simultaneously activating neurons for ”depiction of human figures” (human category), ”transportation-related elements in urban settings” (environment category), and ”anatomically impossible structures” (structure category), providing comprehensive characterization of image samples (more examples in Appendix Appendix B. Qualitative Examples of Image-caption Pairs and Activated Neurons).
@@
LanSE’s comprehensive visual pattern vocabulary enables the construction of four diagnostic metrics that systematically characterize different aspects of generated content: prompt match, visual realism, physical plausibility, and content diversity (detailed definition in Appendix Appendix E. Mathematical Definitions of the Four Metrics). These metrics target complementary aspects that various existing metrics fail to distinguish as shown in Figure 4.
@@
To demonstrate LanSE’s advantages for content analysis, we compare our diagnostic capabilities with widely-adopted baseline methods: Fréchet Inception Distance (FID) [jayasumana2024rethinkingfidbetterevaluation], Inception Score (IS) [barratt2018noteinceptionscore], and CLIP score [chen2025contrastivelocalizedlanguageimagepretraining].
Our analysis aligns with CLIP’s systematic biases identified in various previous works [schrodi2025effectstriggermodalitygap]. While both CLIP and our prompt match metric generally detect severe misalignments (empty upper-right regions in Figure 10), CLIP can assign high scores to outputs with obvious semantic mismatches when partial object alignment masks compositional errors, as shown in Figure 9. In contrast, LanSE’s prompt match evaluation is based on visual patterns and can capture subtle semantic mismatches CLIP fails to detect.
@@
In this era of generative AI, many researchers consider LMMs with prompt engineering as a quick and universal solution for diverse problems [Koco__2023, bubeck2023sparksartificialgeneralintelligence, bang2023multitaskmultilingualmultimodalevaluation]. Comparatively, LanSE can detect fine-grained visual patterns that state-of-the-art LMMs miss, while requiring significantly lower computational resources during inference. We compare LanSE’s specialized error detection capabilities against six state-of-the-art vision-language models on identifying physical plausibility errors. Using 3,410 human-annotated images (2,522 with physics violations, 888 without), we tasked each system with detecting anatomical impossibilities, distorted structures, and physics violations.
@@
We are witnessing a profound societal transformation as AI-generated content gradually permeates nearly every domain of human content creation, fundamentally challenging how we understand, verify, and trust the materials we encounter. Research across multiple domains documents serious societal risks from AI-generated content proliferation, including homogenization of human expression [suzuki2023we] and creative diversity [doshi2024generative], vulnerability to data poisoning attacks in medical systems [alber2025medical], model collapse from recursive training on synthetic data [shumailov2024ai, dohmatob2024strongmodelcollapse], contamination of scientific datasets with undetected artificial material [acion2023generative], and erosion of editorial standards in professional publishing [lindsay2023llms]. These documented challenges underscore the critical need for comprehensive content analysis infrastructure that can systematically characterize AI-generated content properties, enabling informed governance and deployment of these materials across high-stakes domains. LanSE provides the systematic content analysis infrastructure this transformed landscape demands, automatically discovering thousands of interpretable visual patterns that enable comprehensive, human-understandable assessment of AI-generated content properties.
@@
LanSE’s systematic pattern analysis transforms how critical institutions—from hospitals to regulatory agencies—understand and govern AI-generated content. In medical diagnostics and financial decision-making, LanSE provides clinicians and analysts with language-grounded explanations of AI predictions, delivering the transparency and accountability these domains require. Regulators can customize LanSE’s pipeline to detect domain-specific risks for educational, publishing, and safety-sensitive materials, enabling precise content filtering with transparent natural language justifications. For machine learning researchers, LanSE enables systematic dataset analysis by quantifying pattern distributions to reveal bias and guide targeted augmentation, while providing early warning signals for model collapse through diversity measurement [10.1093/jcr/ucaf013, kazdan2025collapsethriveperilspromises]. Beyond analysis applications, LanSE enables integration into training pipelines for interpretable model feedback and supports development of pattern-based retrieval systems where users can find content sharing specific visual characteristics.
For the mechanistic interpretability community, LanSE represents a fundamental reorientation of the field’s purpose—demonstrating for the first time that interpretability modules, including sparse autoencoders and transcoders, previously confined to dissecting model circuits and scientific investigations, can be repurposed as practical infrastructure for urgent societal challenges in content verification, safety assessment, and quality control [bereska2024mechanisticinterpretabilityaisafety, rai2025practicalreviewmechanisticinterpretability, sharkey2025openproblemsmechanisticinterpretability]. While the field has pursued increasingly microscopic questions about model understanding and steering [chen2025personavectorsmonitoringcontrolling, turner2024steeringlanguagemodelsactivation, zou2025representationengineeringtopdownapproach, wu2025axbenchsteeringllmssimple], we show these same mathematical frameworks can directly address the content analysis crisis facing billions of users, establishing interpretability methods not as endpoints for understanding AI systems but as foundational tools for governing their outputs.
While LanSE demonstrates versatility across multiple applications, several avenues could strengthen and extend this method. First, expanding the scale and granularity of discovered visual patterns would enhance evaluation precision, as current assessments remain bounded by our pattern vocabulary. Second, although we successfully adapted LanSE to medical imaging through CXR-LanSE, extending the framework to additional modalities, video, 3D content, or scientific imaging, presents promising opportunities. Third, while we currently demonstrate LanSE primarily for content analysis, dataset characterization, and model correlation studies, we envision integrating LanSE directly into model training pipelines, enabling generative models to receive diverse, interpretable feedback on specific visual patterns during optimization. Fourth, developing adaptive pattern vocabularies that evolve with emerging AI capabilities could maintain LanSE’s effectiveness as generative models advance (see Appendix Appendix I. Limitations).
@@
We use a collection of image-caption pairs from MS-COCO [lin2015microsoftcococommonobjects], Flickr8k and Flickr30k [plummer2016flickr30kentitiescollectingregiontophrase], TextCaps [sidorov2020textcapsdatasetimagecaptioning], NoCaps [Agrawal_2019], and MIMIC-CXR [johnson2019mimiccxrjpglargepubliclyavailable] to build LanSE and CXR-LanSE. We augment these datasets with samples from several generative models [meng2022sdeditguidedimagesynthesis, dalle3, FLUX.1-dev, Kolors-diffusers, zhang2023addingconditionalcontroltexttoimage, malik2023stable] by prompting them with captions from the original collections. We also employ human data annotators to identify erroneous images from the dataset for targeted discovery of physical plausibility patterns (detailed in Appendix Appendix D. Method Extention).
@@
To systematically discover interpretable visual patterns for content analysis, we leverage recent advances in mechanistic interpretability [cunningham2023sparseautoencodershighlyinterpretable, bussmann2025learningmultilevelfeaturesmatryoshka, gao2024scalingevaluatingsparseautoencoders, dunefsky2024transcodersinterpretablellmfeature, paulo2025transcodersbeatsparseautoencoders, lindsey2024sparse]. Studies have shown that enforcing sparsity during neural network training produces monosemantic neurons—individual units that respond to specific, interpretable concepts [makhzani2014ksparseautoencoders, farrell2024applyingsparseautoencodersunlearn, elhage2022toymodelssuperposition, bricken2023monosemanticity], motivating us to adapt these techniques for content analysis.
@@
Sparse autoencoders are sparsity-driven interpretability modules that can comprehensively discover a diverse set of general visual patterns through the reconstruction of one model embedding. We develop multiple sparse autoencoders, each comprising a single-layer encoder $\mathbf{z_{i}}=\mathrm{Enc}(\mathbf{e_{i}})=\mathrm{ReLU}(\mathbf{w_{enc}}\cdot\mathbf{e_{i}}+\mathbf{b_{enc}})$ and a single-layer decoder $\mathrm{Dec}(\mathbf{z_{i}})=\mathrm{ReLU}(\mathbf{w_{dec}}\cdot\mathbf{z_{i}}+\mathbf{b_{dec}})$. A top-$k$ activation scheme [gao2024scalingevaluatingsparseautoencoders] is used to encourage sparsity in the latent space: for each $\mathbf{z_{i}}$, we retain only the $k$ highest-magnitude activations and set the rest to zero. The training objective is as follows:
@@
The neurons discovered by these interpretability modules are initially raw and noisy, requiring systematic curation to identify reliably interpretable patterns. To maximize pattern discovery coverage while enabling quality filtering, we train multiple modules on different data subsets. In total, we train 200 sparse autoencoders and 50 transcoders for LanSE, plus 100 sparse autoencoders for CXR-LanSE, with each module containing 15,000 latent neurons and k=32 for the top-k sparsity constraint. From this large pool of candidate neurons, we apply systematic filtering criteria to identify those with consistent, interpretable activations. We adopt hyperparameters similar to recent developments in mechanistic interpretability [sun2025high, minegishi2025rethinkingevaluationsparseautoencoders] and the approach can be scaled with additional compute and data.
@@
After training the interpretability modules, we obtain a collection of latent neurons, each of them potentially interpretable in natural language. We then employ LMMs [claudehaiku, llama3] to analyze and interpret these latent neurons.
For each neuron, we collect a set of samples that trigger strong activations for this neuron and prompt Claude-3.5-Haiku [claudehaiku] (a large multimodal model) to summarize the common characteristics within these samples. The prompt we use to summarize the visual pattern in natural language is given below (more prompts in Appendix Appendix C. LMM Analysis Prompts):
@@
Based on the natural language summarizations of the visual patterns, we categorize these visual patterns into nine groups, $\mathbf{C}$, including five groups designed for semantics: human, animal, object, activity, and environment; two groups for visual realism: style and artifact; and two groups for physical plausibility: distortion and structure (See Appendix Appendix G. The Choice of LanSE Neuron Groups). We use the following prompt to categorize the visual pattern based on their explanations.
@@
These visual pattern analysis procedures can be formulated as below [luo2023promptengineeringlensoptimal], noticing that the summarizations are obtained automatically during the construction of LanSEs, no assistance from LMMs are used in inference.
@@
Modality distillation refers to training single-modality LanSEs, $\mathrm{S^{img}_{\text{c}}}$ and $\mathrm{S^{txt}_{\text{c}}}$, using outputs from $\mathrm{S_{\text{c}}}(\mathbf{x},\mathbf{y})$ [hinton2015distillingknowledgeneuralnetwork]. The LanSEs are initialized with pretrained encoders and tuned with parameter efficient fine-tuneing [xu2023parameterefficientfinetuningmethodspretrained]. The optimization target is to minimize the distances between the outputs of the single-modality LanSEs and $\mathrm{S_{\text{c}}}(\mathbf{x},\mathbf{y})$: